OpenStudy (mindblast3r):

A can of juice is 680mL, on the nutrition facts it only shows how much sugar your get per 240mL. If it only shows 17g sugar per 240mL, how much sugar do you get for the entire 680mL?

OpenStudy (mathstudent55):

This is a proportion problem. Do you know how to set up a proportion?

OpenStudy (mindblast3r):

i think so.

OpenStudy (mindblast3r):

\[\frac{ 240 }{ 17 }=\frac{ 640 }{ x }\]

OpenStudy (mindblast3r):

17 * 640 ÷ 240 = 45grams of sugar?

OpenStudy (mindblast3r):

@mathstudent55

OpenStudy (mathstudent55):

Great job! Just one small correction. The correct amount is 680, not 640. Your method is correct. Just redo it with the correct number.

OpenStudy (mindblast3r):

oh lol! thank you very very much <3, how did you know it was a proportion problem?

OpenStudy (mathstudent55):

The info the problem gives you is a ratio. 17 g of sugar to 240 ml of drink is a ratio. Then you have another ratio containing an unknown. That is: x g of sugar to 680 ml of drink. A proportion is defined as setting two ratios equal to each other.

OpenStudy (mindblast3r):

okay thank you very very much<3

OpenStudy (mathstudent55):

You are very welcome.